OpenStudy (lgbasallote):

can someone show me the solution for \[\int_{-3}^3 \int_0^{\pi / 2} (y + y^2 \cos x) dx dy\] i got zero but the book says 18 >.<

OpenStudy (lgbasallote):

\[\Large \int_{-3}^3 \int_0^{\pi / 2} (y + y^2 \cos x)dxdy\] i just wanna latex :p

sam (.sam.):

\[\int\limits_{-3}^3[ \int\limits_0^{\pi / 2} (y + y^2 \cos x) dx] dy\] integrate the bracket inside first

OpenStudy (lgbasallote):

i believe i got \[\Large \frac{\pi y}{2} - y^2\]

sam (.sam.):

\[\int\limits \left(y+y^2 \cos (x)\right) \, dx\] i get\[ \frac{1}{2} y (2 y+\pi )\]

OpenStudy (lgbasallote):

oh wait...i got \[\large [yx|_0^{\pi /2} - [y^2 \sin x|_0^{\pi /2}\]

sam (.sam.):

yeah just insert the pi/2 into x

OpenStudy (lgbasallote):

*facepalm*

OpenStudy (lgbasallote):

it's supposed to be plus darn =_=

OpenStudy (lgbasallote):

but im still getting 0 >.<

OpenStudy (lgbasallote):

oh wait nevermind...i saw it

sam (.sam.):

\[y^2\int\limits \cos (x) \, dx+\int\limits y \, dx\] \[y^2 \sin (x)+\int\limits y \, dx\] Integrated inside the brackets i get \[[y^2 \sin (x)+x y]_{0}^{pi/2}\]

sam (.sam.):

ok

OpenStudy (lgbasallote):

seems my mistake lied on \[\frac{3^3}{3} - (-\frac{3^3}{3})\] which then becomes positive